Find the measure of an angle such that the difference between its supplement and 2 times its complement is 70°
Found 2 solutions by mananth, algebrahouse.com:
Answer by mananth(16946)

let angle be x
180-x-2(90-x)=70
180-x-180+2x=70
x=70 the angle

Answer by algebrahouse.com(1659)

Find the measure of an angle such that the difference between its supplement and 2 times its complement is 70°

x = the angle
180 - x = its supplement {supplementary angles add up to 180°
90 - x = its complement {complementary angles add up to 90°}

180 - x - 2(90 - x) = 70 {difference between its supplement and 2 times its complement equals 70°}
180 - x - 180 + 2x = 70 {used distributive property}
x = 70 {combined like terms}

70° is the angle
www.algebrahouse.com